OpenStudy (anonymous):

Can anyone evaluate 3^-4

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

i need it in a fraction

OpenStudy (mayankdevnani):

\[\large \bf 3^{-4}\] Always remember this property:- \[\large \bf x^{-y}=\frac{1}{x^{y}}\] so,we get \[\large \bf \frac{1}{3^4}=\frac{1}{81}\]

OpenStudy (mayankdevnani):

this is your answer ^^^ @CoolRanchDoritos

OpenStudy (ibbutibbu.):

oh yeah ur right

OpenStudy (mayankdevnani):

understood? @CoolRanchDoritos

OpenStudy (anonymous):

thank you

OpenStudy (mayankdevnani):

welcome :)